OpenStudy (anonymous):

Find the first three terms of the geometric sequence where the first term is 4 and the common ratio is 1/2 HELP MEDALS GIVEN

OpenStudy (aum):

Multiply first term by common ratio to get the second term. Multiply second term by common ratio to get the third term.

OpenStudy (anonymous):

HOW PLEASE can you shoiw me

OpenStudy (aum):

First term is 4 Second term = First term * common ratio = 4 * 1/2 = ?

OpenStudy (anonymous):

2 right

OpenStudy (aum):

Yes. First term is 4 Second term = First term * common ratio = 4 * 1/2 = 2 Third term = Second term * 1/2 = 2 * 1/2 = 1 So the first three terms of the geometric sequence are: 4, 2, 1

OpenStudy (anonymous):

i have to pass this test i ahve the answer choices

OpenStudy (aum):

Is 4, 2, 1 among the choices?

OpenStudy (anonymous):

omgggg thank can you help with some more pleaseee

OpenStudy (aum):

I am not supposed to be helping with tests. Sorry.

OpenStudy (anonymous):

this is the study guide i have to know the material to pass the test this isnt the actually test thats why im asking for help

OpenStudy (aum):

I can assist you but you have to do the work yourself. I will try one more.

OpenStudy (anonymous):

Find the first three terms of the sequence: dn=3/5n-1

OpenStudy (aum):

Is it \[ d_n = \frac{3}{5n-1}\] ?

OpenStudy (anonymous):

yesss can you please whow me how you typed that???

OpenStudy (aum):

You can click on the Equation button below the reply box and choose what you want from the right hand side.

OpenStudy (anonymous):

do you know how to put it in wolfram alpha ?

OpenStudy (aum):

It is so simple to do it than to come up with instructions for wolfram to do it.

OpenStudy (aum):

Just put n = 1 on the right hand side and evaluate. That will be the first term. Put n = 2 on the right hand side and evaluate. That will be the second term.

OpenStudy (anonymous):

okay lollll im a little slow with things

OpenStudy (aum):

3 / (5n - 1) n = 1 3 / (5 - 1) = 3/4 is the first term. put n = 2 in 3 / (5n - 1) and evaluate to get the second term. put n = 3 in 3 / (5n - 1) and evaluate to get the third term.

OpenStudy (anonymous):

okay so the answer is

OpenStudy (anonymous):

nevermid i got ittttt

OpenStudy (anonymous):

hey how do you put this in a calculator on wolfram

OpenStudy (anonymous):

a6= 4n^2+5/7n^3+n